Little Dome C - Beyond EPICA Drilling Site (75.29917 degS, 122.44516 degE)

Weekly REPORT - 14-21 November 2021

·       The maintenance of the two PBs and of the caterpiller has been completed. New radios were also installed on the vehicles

·       First raid to LDC Wednesday 17, with PB and one fuel tank. Big accumulation of snow at the camp, much more than expected. The tents are in excellent condition, but the camp needs to be cleaned before starting the various planned activities.

·        The strong wind on Thursday and Friday, with temperatures below -60°C, made it impossible to carry out outdoor activities. The ONAN power generator, inside ITASE module, has been turned on and left in that way.

·        Second raid to LDC on Saturday 20, with PB + fuel tank and the other PB, to start cleaning the camp. The area for the ITASE modules has been cleaned and now is ready for them.


Saverio Panichi